7. Consider this simplified sequence: A farmer starts with a unit of wheat and sells it to a

miller for 20 cents. The miller grinds it into flour and sells it to a baker for 50 cents. The

baker converts all the flour into a loaf of bread and sells it to a grocer for $2.00. The grocer

then sells that bread to a customer for $3.35.

a) What is the total value added by the grocer?

b) What is the total value added by all these stages of production?

c) How much does GDP increase?


1
Expert's answer
2021-06-29T10:25:46-0400

a)


"3.35-2=1.35"

b)


"3.35-0.2=3.15"

c)


"0.2+0.5+2+3.35=6.05"